# Value Residual Learning For Alleviating Attention Concentration In Transformers

### v3

@@ 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
Compared to previous methods, SVFormer is the first method that decouple value vectors from attention. Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers. However, the results in Fig. 5 show that sharing values has less negative impact compared with sharing keys.
Figure 5: Ablation study of sharing different parts of attention every two layers.
4 Pretrain Experiments
@@ 4.1.1 Training Details
Following Brandon et al. (2024), we choose the Llama-like architecture and SlimPajama (Soboleva et al., 2023) data for main experiments. Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout. For slimpajama, we randomly sample nearly 20B tokens according to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base”. The details of training data can be found in Table 2 in Appendix.
@@
Unless otherwise noted, we train all models using AdamW optimizer with 0.1 weight decay, $\beta_{1}=0.9$, $\beta_{2}=0.95$ and the max grad norm 1.0. The batch size is set to be around 2M tokens (Zhang et al., 2024) with a sequence length of 2,048 and the total steps is fixed 10,000 steps (Kaplan et al., 2020). We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively. The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Zhou et al., 2024, Wei et al., 2023). The detailed hyperparameters for models of various sizes and different training sequence lengths in Appendix A.5. Moreover, All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16. We adopt deepspeed zero-2 optimizer and flash attention mechanism.
4.1.2 Relative Training Loss Curve On SlimPajama
We trained all models for only one epoch on SlimPajama subsets, and primarily use training loss to compare different models. Furthermore, we use the relative training loss curve for better visualizing the difference among different models from the perspective of loss landscape. Specifically, for each method, we will subtract the smoothed training curve of the vanilla Transformer, obtained under the same experimental settings, from the smoothed training curves of the method. The smoothing is done using a window size of 10 steps or 100 steps.
4.1.3 Entropy For Analyzing Attention Concentration Effects
Given the attention matrix $\mathbf{A}\in\mathbb{R}^{l\times l}$ at one layer, we use entropy $\bm{e}$ to represent its concentration effect. To obtain entropy $\bm{E}$, calculate the importance vector $\mathbf{a}=\frac{1}{l}\sum_{j=1}^{l}A_{ij}$ firstly where $\mathbf{A}$ is a lower triangular matrix. The entropy can be formulated as follows:
$\bm{e}$ $=-\sum_{i=1}^{l}\bm{a}_{i}^{\prime}\log{\bm{a}^{\prime}_{i}}.$ (11)
where $a_{i}^{{\prime}_{i}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
4.1.4 Spectral Decomposition For Analyzing Representations
Spectral Decomposition is a classical method to analyze the representations of models. Zhu et al. (2021) suggests that the eigenvectors with larger eigenvalues are more transferable. Here we use spectral decomposition to analyze the feature space of value $\bm{v}$ of one layer as following:
$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{j=1}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.\vspace{-3mm}$ (12)
where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $j=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
@@ 4.1.4 Spectral Decomposition For Analyzing Representations
Figure 6: (Left) Validation loss as model size scales from 82M to 468M parameters. (Right) Validation loss for the 468M parameter model evaluated every 2B tokens. ResFormer achieves approximately 10.4%-13.6% reduction in both model parameters and training data.
Figure 7: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths. (Right) Relative training loss of various Transformer variants compared to the vanilla Transformer model, with model size fixed at 82M parameters.
4.2 ResFormer vs. vanilla Transformer
We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths. In Fig. 7 (Left), ResFormer consistently outperforms vanilla Transformer throughout training across different training sequence lengths. Additionally, the results in Fig. 7 (Right) illustrate that ResFormer outperforms DenseFormer and NeuTRENO. Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.

### v4

@@ 3.4 SVFormer
Beyond ResFormer, SVFormer adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through $\mathbf{U}_{n}=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers. However, the results in Table  1 show that sharing values has less negative impact compared with sharing keys.
4 Experiments
@@ 4.1 Setting
Following Brandon et al. (2024), we choose the Llama-like architecture and SlimPajama (Soboleva et al., 2023) data for main experiments. Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout. For slimpajama, we randomly sample nearly 20B tokens based to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base". See Table 12 in Appendix for data details.
@@
Unless otherwise noted, we train all models using AdamW optimizer with 0.1 weight decay, $\beta_{1}=0.9$, $\beta_{2}=0.95$ and the max grad norm 1.0. The batch size is set to be around 2M tokens (Zhang et al., 2024) with a sequence length of 2,048 and the total steps is fixed 10,000 steps (Kaplan et al., 2020). We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively. The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Zhou et al., 2024; Wei et al., 2023). The detailed hyperparameters for models of various sizes and different training sequence lengths in Appendix A.4. Moreover, All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16. We adopt deepspeed zero-2 optimizer and flash attention mechanism.
Evaluation
For model evaluation and comparison, we primarily utilized the average validation loss across seven domains, computed on the entire SlimPajama validation split. Additionally, we randomly selected a fixed set of 1,000 sample sequences for subsequent visualization analysis.
4.2 Scaling Analysis of ResFormer
@@
Figure 2: (Left) Validation loss as model size scales from 82M to 468M parameters. (Right) Validation loss for the 468M parameter model evaluated every 2B tokens. ResFormer achieves approximately 13.3%-15.4% reduction in both model parameters and training data.